To pay off a loan of 7000 due now and a loan of 2000 in 14 month's time olarato agrees to take three payments in two, 5 and 10 months time respectively. The second payment is to be double the first and the third payment is to be triple the 1st. what is the size of the payment at month 5 if interest is calculated at 16% per year compounded monthly


1
Expert's answer
2022-03-07T17:26:04-0500

Total loan amount:

7000+2000=9000


find the monthly payment:

"r=\frac{16}{12}=1.33"


"A=9000(0.0133+\frac{0.0133}{(1+0.0133)^{14}})=219.74"

A1=219.74

"A2=2\times219.74=439.47"

"A3=3\times219.74=659.21"

"A5=5\times219.74=1098.67"